# Space-time first-order correlations of an open Bose Hubbard model with incoherent pump and loss

### v1

@@ 5.2 Position of the excitation spectral lines for small hopping
We use the one-site model, i.e. we set $J=0$. In this case, the Hamiltonian becomes local in position space
@@ 5.3 Spectral function for the ground state of the Bose-Hubbard model
Figure 6: Spectral function $A(k,\omega)$ for the ground state of the Bose-Hubbard model under unitary evolution (negative values in grey). Left panel: Exact solution without truncation at finite interaction ($L=12$, $N=2$ particles, $U/J=10$). Right panel: larger system size ($L=24$, $N=6$, $U/J=10$, with $N_{s}=3$ obtained with DMRG).
Let us comment on the results for the closed system. In the lowest-energy excitation manifold, we clearly identify the Lieb-I and Lieb-II branches, as well as the third branch predicted for the lattice case[17]. The predictions for the dispersion branches in the infinite interaction limit is also shown in Fig. 6. Notice that in this limit all the branches are slightly shifted upwards with respect to the finite-$U$ spectrum, since they are centered at $\omega=\mu-2J$ and the value of the chemical potential $\mu$ at infinite interactions is larger than the one at finite interactions. At frequencies $\omega\simeq\mu+U+2J$, we also see in the figure a dispersive doublon branch.
@@ 5.4 Comparison of unitary evolution and Lindblad evolution of the NESS
Figure 7: Spectral function $A(\omega,k)$ of the Markovian open system dynamics in the NESS. Left: Under unitary dynamics; Right: Under Lindbladian dynamics. The parameters are: $U=500,\,J=50,\,z=0.2,\,L=8,\,N_{s}=3$. The horizontal white lines are as in Fig. 5.
@@ 6 Conclusions and perspectives
The study of the bosonic quantum gas coupled to an environment is a present problem of particular interest. Open quantum systems play a crucial role in the current noisy intermediate-scale quantum (NISQ) era [66], with quantum science and quantum engineering bringing forth a variety of platforms for quantum simulations of emergent collective many-body phenomena. Notwithstanding the exceptional technological advances, photonic platforms remain intrinsically dissipative and we foresee as particularly important for the near future the characterisation of the interplay of the open-system dynamics of these setups with the coherent Hamiltonian dynamics, which will be characterised by increasingly strong Kerr non-linearities.
@@
In this article we have studied the space-time behaviour of two-point correlation functions and the spectral function for a driven-dissipative system of interacting bosons on a lattice, subjected to uniform incoherent pump and one-body losses. Our starting point has been the analytical expression for the NESS density matrix of the system derived in Ref. [58]. We have used several techniques, namely numerical exact diagonalisation, Keldysh formalism and strong coupling approaches, to provide a complete understanding of the excitation spectrum on top of the non-equilibrium steady state. We have found that the decay of the retarded Green’s function is exponential in time, and the decay rate is renormalised by interactions and increases at both increasing interactions and tunnel amplitude. The spectral function displays at low frequency a quadratic-like branch, very different from the linear branch found for the excitations on top of the ground state of the closed system, for which we have calculated the spectral function at finite interactions. At large interactions, we have identified, both in the open and in the closed system, a dispersive doublon branch at energy of order $U$. By performing calculations for various system sizes in the closed system case, we have shown that specific features of the spectrum clearly emerging at large system sizes are also recognisable in smaller systems, of typical size accessible via numerical diagonalisation in the open case.
@@
In outlook, on a technical level, it would be interesting to push the calculations for the spectral function of the open system to larger system sizes using techniques based on matrix-product states such as those detailed in Refs. [67, 68]. This would provide conclusive evidence about the presence of excitations at negative energies (the ‘ghost’ branches) which, according to Bogoliubov theory, are predicted to be more elusive in the driven-dissipative case than in the closed one [69, 70, 71] and could not be resolved in our current numerical calculations. On a fundamental level, our work opens very interesting perspectives for the possibility of observing the effects of interactions in out-of-equilibrium systems within actual experiments, where short one-dimensional lattices of small photonic resonators can be engineered.
@@ Acknowledgements
We warmly thank I. Carusotto, L. Garbe and S. Diehl for fruitful discussions, and L. Rosso for significant help in the early stages of the project. MZ thanks for visiting the LPTMS and using their computational infrastructure.
@@
AM acknowledges funding from the Quantum-SOPHA project ANR-21-CE47-0009; This work is part of HQI (www.hqi.fr) initiative and is supported by France 2030 under the French National Research Agency grant number ANR-22-PNCQ-0002; LM acknowledges funding from the ANR project LOQUST ANR-23-CE47-0006-02; LC acknowledges support from Institut Universitaire de France (IUF).
Appendix A Detailed calculations
@@ A.1 Bosonic NESS density matrix
We show in this section that the ansatz (16) is a NESS of the Lindblad evolution (2). The Hamiltonian conserves the particle number and hence commutes with each identity operator in each particle number subspace. Furthermore, the non-hermitian Hamiltonian terms in (2) can be written as
@@
$\sum_{i}b_{i}^{\dagger}\rho_{\text{NESS}}b_{i}$ $=\sum_{i}\sum_{N}\omega_{N-1,L}\sum_{\begin{subarray}{c}\{l_{r}\}_{r=1,...,L;N}\\ \{n_{r}\}_{r=1,...,L;N}\end{subarray}}|\{l_{p}\}\rangle\langle\{n_{s}\}|\sum_{\{m_{r}\}_{r=1,...,L;N-1}}\langle\{l_{p}\}|b_{i}^{\dagger}|\{m_{r}\}\rangle\langle\{m_{r}\}|b_{i}|\{n_{s}\}\rangle$
$=\sum_{N}\omega_{N-1,L}\sum_{\begin{subarray}{c}\{l_{r}\}_{r=1,...,L;N}\\ \{n_{r}\}_{r=1,...,L;N}\end{subarray}}|\{l_{p}\}\rangle\langle\{n_{s}\}|\langle\{l_{p}\}|\sum_{i}b_{i}^{\dagger}b_{i}|\{n_{s}\}\rangle$
$=\sum_{N}N\omega_{N-1,L}\mathbb{1}_{N}\;,$ (30)
@@
$\sum_{i}b_{i}\rho_{\text{NESS}}b_{i}^{\dagger}$ $=\sum_{N}(N+L)\omega_{N+1,L}\mathbb{1}_{N}\;.$ (31)
@@
$\mathcal{L}_{\text{loss}}[\rho_{\text{NESS}}]+\mathcal{L}_{\text{gain}}[\rho_{\text{NESS}}]$ $=\sum_{N}\mathbb{1}_{N}\Big{[}\gamma_{l}\Big{\{}(N+L)\omega_{N+1,L}-N\omega_{N,L}\Big{\}}$
$\qquad\qquad+\gamma_{p}\Big{\{}N\omega_{N-1,L}-(N+L)\omega_{N,L}\Big{\}}\Big{]}$
$=\sum_{N}\mathbb{1}_{N}\omega_{N-1,L}\Big{[}\gamma_{l}\Big{\{}z^{2}(N+L)-zN\Big{\}}+\gamma_{p}\Big{\{}N-z(N+L)\Big{\}}\Big{]}$
@@ A.3 Equal-time correlation functions
The geometrical series in the second line only converges if $|z|<1$. All together, we find the result (19) as expected from the calculation of the total particle number in the NESS.
@@ Appendix B Normalisation with truncation of the local Hilbert space
This section gives the normalisation of the NESS density matrix in each $N$-particle sector depending on the system size $L$ and on the local Hilbert space dimension $d_{l}$. For this, one has to solve a restricted star (balls/particles) and bar (box separations) problem which is well known in combinatorics
@@
and $x_{i}\in{0,...,N}$. The latter condition is equivalent to the existence of an upper bound for the integers $x_{i}<d_{l}$$\forall i$. This problem can be solved by using the inclusion-exclusion principle on the lower-bound integer sum. The latter is given by the previous sum with the constraint $x_{i}\geq a_{i}$. We substitute $x_{i}^{\prime}\coloneqq x_{i}-a_{i}$, such that the modified problem is the unbounded one with $x_{i}^{\prime}\geq 0$:
@@
then we subtract all the cases where there is at least one box with $x_{i}\geq d_{l}$ (solving the lower-bound integer sum):
$\begin{pmatrix}L\\ 1\end{pmatrix}\begin{pmatrix}L+N-1-d_{l}\\ L-1\end{pmatrix}\;,$ (48)
and finally add the two set intersections, such that
$D^{d_{l}}_{L,N}\coloneqq\#\text{states}(d_{l},L,N)=\sum_{k=0}^{k(d_{l}-1)\leq N}(-1)^{k}\begin{pmatrix}L\\ k\end{pmatrix}\begin{pmatrix}L+N-1-kd_{l}\\ N-1\end{pmatrix}\;.$ (49)
Appendix C Steady state solution for fermions
@@ C.1 NESS density matrix
In finite dimensional Hilbert spaces, there always exists a steady state solution of the Lindbladian superoperator [43]. Starting from the same ansatz  (16) as in the main text, we obtain from the normalisation of the state that
$\mathcal{N}^{F}_{L}=\frac{1}{(1+\gamma)^{L}}\;,$ (50)
where we used that the dimension of the Hilbert space for $N$ fermions in a system of length $L$ corresponds to $D^{F}_{L,N}=\binom{L}{N}$. This is exactly the size of the unit matrix $\operatorname{Tr}\mathbb{1}_{N}=D^{F}_{L,N}$. This provides the NESS density matrix for fermions.
@@ C.2 Equal-time two-point correlation functions
$\sum_{\{m_{r}\}_{r=1,...,N}}\langle{\{m_{r}\}}|c_{i}^{\dagger}c_{j}|{\{m_{r}\}}\rangle$ $=\sum_{i,\{m_{r}\}_{r=1,...,N-1}\in\{1,...,L\}/\{i\}}\langle{\{m_{r}\}}|c_{i}^{\dagger}c_{j}|{\{m_{r}\}}\rangle\delta_{i,j}$

### v2

@@ 6.2 Position of the excitation spectral lines for small hopping
We use the single site model, i.e. we set $J=0$. In this case, the Hamiltonian becomes local in position space
@@ 6.3 Spectral function for the ground state of the Bose-Hubbard model
Figure 6: Spectral function $A(k,\omega)$ for the ground state of the Bose-Hubbard model under unitary evolution (negative values in grey). The blue and red lines indicate the Lieb-I and II branches respectively, the green branch is a feature of the lattice model, see text. Left panel: Exact diagonalisation solution without truncation in the local Hilbert space at finite interaction ($L=12$, $N=2$ particles, $U/J=10$), compare with Appendix G.5. Right panel: Larger system size ($L=24$, $N=6$, $U/J=10$, with $N_{s}=3$ obtained with DMRG).
Let us comment on the results for the closed system. In the lowest-energy excitation manifold, we clearly identify the Lieb-I and Lieb-II branches, as well as the third branch predicted for the lattice case[17]. The predictions for the dispersion branches in the infinite interaction limit is also shown in Fig. 6. Notice that in this limit all the branches are slightly shifted upwards with respect to the finite-$U$ spectrum, since they are centered at $\omega=\mu-2J$ and the value of the chemical potential $\mu\,=\pi^{2}J(\frac{N}{L})^{2}$ at infinite interactions is larger than the one at finite interactions. At frequencies $\omega\simeq\mu+U+2J$, we also see in the figure a dispersive doublon branch.
@@
6.4 Spectral function of the finite-temperature equilibrium state
In this section, we calculate the spectral function of an equilibrium state at finite temperature, and we highlight the differences with the spectral function of the open quantum system.
6.4.1 Equilibrium spectral function
The spectral function is obtained by performing thermal averages (see again Eq. (7)) of the two-point correlations functions, assuming a grand-canonical density matrix
$\rho_{GC}(\beta,\mu)=\frac{1}{Z}\text{e}^{-\beta(H-\mu N)}\,\;,$ (30)
with $Z={\rm Tr}[\text{e}^{-\beta(H-\mu N)}\,]$. In order to compare the results with the spectral functions obtained in the open system, we choose the same size and average filling of the system in the two cases. We obtain the chemical potential $\mu$ by numerically calculating the average occupation $\langle N\rangle(\mu)$ and inverting the relation, as illustrated in Fig. 7.
Figure 7: Average filling $\langle N\rangle$ of the system depending on the chemical potential $\mu$ given a grand-canonical distribution (30), with system parameters $U/J=10\,\,L=8,\,N_{s}=3$. Left panel: Low-temperature regime, with $\beta=100J^{-1}$. The Mott lobes correspond to the plateaus. Right panel: High-temperature $\beta=0.01J^{-1}$.
The spectral function is shown in Fig. 8 for two values of the temperature. In the extremely low-temperature regime (left panel), the spectral function is very close to the ground-state results presented in Fig. 6, as expected. In the extremely high-temperature regime (right panel) the spectral lines are blurred with a broadening which, in the upper branch, reaches the whole bandwidth. We stress that the linewidth of the high-temperature case is different from the case of a NESS, shown in Fig. 5, although corresponding to the same values of interaction strength and filling. This shows that although the NESS formally resembles an infinite-temperature state, the time-dependent observables on top of this state are different. Note, that the system is not at infinite temperature, as out-of-equilibrium there is no temperature. Moreover, there is an additional degree of freedom, which is the pump to loss ratio $z$. The comparison of Fig. 5 (right panel) to Fig. 9 (left panel) shows that, by choosing a different ratio of the coherent to the dispersive parameters –while keeping the filling and the ratio $J/U$ constant– one can tune the linewidth, which is not possible in equilibrium.
A more general argument for showing that the NESS (9) is different from a thermal-equilibrium state is obtained by examining the action (4.1): the latter reveals an explicit breaking of the time-reversal symmetry [70] which defines systems at equilibrium (see further details in Appendix E.5). This symmetry, implying an infinite hierarchy of fluctuation-dissipation relations, does not hold for the out-of-equilibrium system. This explicitly implies a different number of degrees of freedom in and out-of-equilibrium. In equilibrium, a semi-classical (high temperature) non-relativistic bosonic theory has a term in the action of the form $\varphi_{q}^{*}\text{i}(\partial_{t}-\kappa)\varphi_{c}+c.c.+4\text{i}\kappa T|\varphi_{q}|^{2}$, which at a given temperature $T$ imposes a fixed relation between the noise and the loss term (sharing the same $\kappa$). In the non-equilibrium action (4.1), these two terms are no longer related by a temperature, but they can vary independently, and they undergo different renormalisations.
Figure 8: Spectral function within an equilibrium grand-canonical state (30) at strong interactions $U/J=10$. Left panel: Low-temperature regime $\beta=100J^{-1}$ (comparison with exact-diagonalisation solution as in Fig. 6). Right panel: High-temperature regime $\beta=0.01J^{-1}$. The average particle number is set to $\langle N\rangle=2$ in order to have the same filling as in the open-system case. The other parameters used are $L=8$, $N_{s}=3$.
@@ 6.5 Comparison of unitary evolution and Lindblad evolution of the NESS
Figure 9: Spectral function $A(\omega,k)$ of the temporal evolution in the NESS. Left panel: Under unitary dynamics (see Appendix G.5); Right panel: Under Lindbladian dynamics. The parameters are: $U=500\gamma_{l},\,J=50\gamma_{l},\,z=0.2,\,L=8,\,N_{s}=3$. The horizontal white lines are as in Fig. 5.
@@ 7 Conclusions and perspectives
In this article we have studied the space-time behaviour of two-point correlation functions and the spectral function for a driven-dissipative system of interacting bosons on a lattice, subjected to uniform incoherent pump and one-body losses. Our starting point has been the analytical expression for the NESS density matrix of the system derived in Ref. [68]. We have used several techniques, namely numerical exact diagonalisation, Keldysh formalism and strong coupling approaches, to provide a complete understanding of the excitation spectrum on top of the non-equilibrium steady state. The numerical results that we have obtained cover the fully quantum evolution, allowing us to access in particular the strongly interacting, short distance regime that cannot be reached by field theoretical methods relying on mean-field or perturbative schemes. Our numerical results are confirmed by various exact results in the closed, as well as zero-dimensional system.
We have found that the decay of the retarded Green’s function is exponential in time, and the decay rate is renormalised by interactions and increases at both increasing interactions and tunnel amplitude. The spectral function displays at low frequency a quadratic-like branch, very different from the linear branch found for the excitations on top of the ground state of the closed system and for excitations on top of the equilibrium system at low temperature for which we have calculated the spectral function at finite interactions. At large interactions, we have identified, both in the open and in the closed system, a dispersive doublon branch at energy of order $U$. By performing calculations for various system sizes in the closed system case, we have shown that specific features of the spectrum clearly emerging at large system sizes are also recognisable in smaller systems, of typical size accessible via numerical diagonalisation in the open case. By calculating the spectral function for an equilibrium state at finite temperature, we have also highlighted the difference with respect to the out-of-equilibrium one, where, in particular, the linewidth can be tuned by changing the ratio of drive to dissipation rate.
@@
In outlook, on a technical level, it would be interesting to develop new methods to push the calculations for the spectral function of the open system to larger system sizes. This would provide conclusive evidence about the presence of excitations at negative energies (the ‘ghost’ branches) which, according to Bogoliubov theory, are predicted to be more elusive in the driven-dissipative case than in the closed one [75, 76, 77] and could not be resolved in our current numerical calculations. On a fundamental level, our work opens very interesting perspectives for the possibility of observing the effects of interactions in out-of-equilibrium systems within actual experiments both in condensed-matter and photonic platforms, where short one-dimensional lattices of small photonic resonators can be engineered, and in ultracold atoms in optical lattices, where drive and dissipation are adjustable.
@@ Acknowledgements
We warmly thank I. Carusotto, L. Garbe and S. Diehl for fruitful discussions, and L. Rosso for participating to the early stage of this work. MZ thanks for visiting the LPTMS and using their computational infrastructure.
@@
AM acknowledges funding from the Quantum-SOPHA project ANR-21-CE47-0009; This work is part of HQI (www.hqi.fr) initiative and is supported by France 2030 under the French National Research Agency grant number ANR-22-PNCQ-0002; LM acknowledges funding from the ANR project LOQUST ANR-23-CE47-0006-02; LC and LM acknowledge support from Institut Universitaire de France (IUF).
Appendix A Detailed characterisation of the non-equilibrium steady state
@@ A.1 Bosonic NESS density matrix
We show in this section that the Ansatz (9) is a NESS of the Lindblad evolution (2). We abbreviate the weight in each particle sector of the density matrix as $\omega_{N}=\frac{z^{N}}{\mathcal{N}}$. The Hamiltonian conserves the particle number and hence commutes with each identity operator in each particle number subspace. Furthermore, the non-hermitian Hamiltonian terms in (2) can be written as
@@
$\sum_{i}b_{i}^{\dagger}\rho_{\text{NESS}}b_{i}$ $=\sum_{i}\sum_{N}\omega_{N-1}\sum_{\begin{subarray}{c}\{l_{r}\}_{r=1,...,L;N}\\ \{n_{r}\}_{r=1,...,L;N}\end{subarray}}|\{l_{p}\}\rangle\langle\{n_{s}\}|\sum_{\{m_{r}\}_{r=1,...,L;N-1}}\langle\{l_{p}\}|b_{i}^{\dagger}|\{m_{r}\}\rangle\langle\{m_{r}\}|b_{i}|\{n_{s}\}\rangle$
$=\sum_{N}\omega_{N-1}\sum_{\begin{subarray}{c}\{l_{r}\}_{r=1,...,L;N}\\ \{n_{r}\}_{r=1,...,L;N}\end{subarray}}|\{l_{p}\}\rangle\langle\{n_{s}\}|\langle\{l_{p}\}|\sum_{i}b_{i}^{\dagger}b_{i}|\{n_{s}\}\rangle$
$=\sum_{N}N\omega_{N-1}\mathbb{1}_{N}\;,$ (33)
@@
$\sum_{i}b_{i}\rho_{\text{NESS}}b_{i}^{\dagger}$ $=\sum_{N}(N+L)\omega_{N+1}\mathbb{1}_{N}\;.$ (34)
@@
$\mathcal{L}_{\text{loss}}[\rho_{\text{NESS}}]+\mathcal{L}_{\text{gain}}[\rho_{\text{NESS}}]$ $=\sum_{N}\mathbb{1}_{N}\Big{[}\gamma_{l}\Big{\{}(N+L)\omega_{N+1}-N\omega_{N}\Big{\}}$
$\qquad\qquad+\gamma_{p}\Big{\{}N\omega_{N-1}-(N+L)\omega_{N}\Big{\}}\Big{]}$
$=\sum_{N}\mathbb{1}_{N}\omega_{N-1}\Big{[}\gamma_{l}\Big{\{}z^{2}(N+L)-zN\Big{\}}+\gamma_{p}\Big{\{}N-z(N+L)\Big{\}}\Big{]}$
@@ A.3 Equal-time correlation functions
The geometrical series in the second line only converges if $|z|<1$. Altogether, we find the result (12) as expected from the calculation of the total particle number in the NESS.
@@ Appendix B Normalisation with truncation of the local Hilbert space
This section gives the normalisation of the NESS density matrix in each $N$-particle sector depending on the system size $L$ and on the local Hilbert space dimension $N_{s}$. For this, one has to solve a restricted star (balls/particles) and bar (box separations) problem which is well known in combinatorics: How many possible sets $\{x_{i}\}_{i\in\{1,\dots,L\}}$ do exist, such that
@@
with $x_{i}\in\{0,...,N\}$?
The local Hilbert space dimension $N_{s}$ is equivalent to the existence of an upper bound for the integers $x_{i}<N_{s}$, $\forall i$. Fermions have an upper bound of $N_{s}=2$, bosons have no upper bound, but in order to get the correct numerical weight of the state, we need to calculate this problem with an arbitrary upper bound $N_{s}$. This problem can be solved by using the inclusion-exclusion principle on the lower-bound integer sum. The latter is given by the previous sum with the constraint $x_{i}\geq a_{i}$. We substitute $x_{i}^{\prime}\coloneqq x_{i}-a_{i}$, such that the modified problem is the unbounded one with $x_{i}^{\prime}\geq 0$:
@@
then we subtract
$\begin{pmatrix}L\\ 1\end{pmatrix}\begin{pmatrix}L+N-1-N_{s}\\ L-1\end{pmatrix}\;,$ (51)
which corresponds to the number of cases where there is at least one box with $x_{i}\geq N_{s}$. In the same manner, we construct all cases of at least $k$ boxes with $x_{i}\geq N_{s}$ and by adding the two set intersections, we obtain the required result. The number of possible sets to have in total $N$ identical particles on $L$ distinguishable sites with maximally $N_{s}-1$ particles per site is
$D^{N_{s}}_{L,N}\coloneqq\#\text{states}(N_{s},L,N)=\sum_{k=0}^{k(N_{s}-1)\leq N}(-1)^{k}\begin{pmatrix}L\\ k\end{pmatrix}\begin{pmatrix}L+N-1-kN_{s}\\ N-1\end{pmatrix}\;.$ (52)
For the special cases of fermions and spin-$\frac{1}{2}$ with $N_{s}=2$ and (not truncated) bosons $N_{s}=\infty$, this simplifies (compare App. C and D to (50).
Appendix C Steady-state solution for fermions
@@ C.1 NESS density matrix
In finite dimensional Hilbert spaces, there always exists a steady-state solution of the Lindbladian superoperator [52]. Starting from the same Ansatz (9) as in the main text, we obtain from the normalisation of the state that
$\mathcal{N}^{F}_{L}=\frac{1}{(1+z)^{L}}\;,\quad\qquad z=\frac{\gamma_{p}}{\gamma_{l}}\in[0,\infty)\;,$ (53)
where we used that the dimension of the Hilbert space for $N$ fermions in a system of length $L$ corresponds to $D^{F}_{L,N}=\binom{L}{N}$. This is exactly the size as the unit matrix $\operatorname{Tr}\mathbb{1}_{N}=D^{F}_{L,N}$. This provides the NESS density matrix for fermions.
@@ C.2 Equal-time two-point correlation functions
$\sum_{\{m_{r}\}_{r=1,...,N}}\langle{\{m_{r}\}}|c_{i}^{\dagger}c_{j}|{\{m_{r}\}}\rangle$ $=\sum_{i,\{m_{r}\}_{r=1,...,N-1}\in\{1,...,L\}\/\{i\}}\langle{\{m_{r}\}}|c_{i}^{\dagger}c_{j}|{\{m_{r}\}}\rangle\delta_{i,j}$